# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

### v1

@@ 3 Bijections between strongly restricted permutations and restricted combinations
A strongly restricted permutation $\pi$ of the set $\mathbb{N}_{n}$ is a permutation for which the number of permissible values of $\pi(i)-i$ for each $i\in\mathbb{N}_{n}$ is less than a finite number independent of $n$ [17]. Here we give two bijections between restricted combinations and types of strongly restricted permutations where the permutations satisfy $\pi(i)-i\in\mathcal{D}$. Our first bijection concerns a strongly restricted permutation where $\mathcal{D}=\{-1,0,-1\}$ and any $m$ consecutive $\pi(i)$ cannot differ from one another by more than $m$.
@@
The permutation $\pi$ of $\mathbb{N}_{n+1}$ corresponding to a subset $\mathcal{S}$ of $\mathbb{N}_{n}$ is formed as follows. If $i\in\mathcal{S}$ (where $i=1,\ldots,n$) then $\pi(i)=i+1$ and $\pi(i+1)=i$ (i.e., a pair of exchanged digits) whereas if neither $i$ nor $i-1$ are in $\mathcal{S}$ (for $i=2,\ldots,n+1$) then $\pi(i)=i$. As $\mathcal{Q}$ contains 1, the pairs of exchanged digits in the permutation cannot overlap. Given that condition (a) is satisfied, the only way for (b) to be violated would be if $\pi(i+1)=i$ and $\pi(i+m)=i+m+1$ but this could only occur if both $i$ and $i+m$ were in $\mathcal{S}$ which is impossible since $m\in\mathcal{Q}$. The process of forming a permutation from a subset is clearly reversible and so the bijection is established. ∎
Theorem 7 has connections to the following sequences in the OEIS [20]: A000930 ($S^{\{1,2\}}_{n}$) and A102547 ($S^{\{1,2\}}_{n,k}$), its associated triangle; A130137 ($S^{\{1,3\}}_{n}$), which we mention again in the section on bit strings; A263710 ($S^{\{1,4\}}_{n}$), which was the sequence that led to the discovery of the bijection and we will now prove the conjectured generating function for that entry; and A374737 ($S^{\{1,5\}}_{n}$), which was recently added by the author.
@@
The generating function $G(x)$ for the number of restricted combinations of $\mathbb{N}_{n}$ with disallowed differences $\{1,4\}$ is given by
$G(x)=\frac{1+x+x^{2}+x^{3}+2x^{4}+x^{6}}{1-x-x^{3}+x^{4}-2x^{5}+x^{6}-x^{7}}.$ (13)
@@
The $\mathcal{Q}$-comb is a $(2,2,1)$-comb. The metatile-generating digraph for tiling with squares and such combs (Fig. 2) has two inner cycles ($SC_{[3]}$ and $C_{[2]}C_{[3]}$) which have lengths $L_{1}=3$ and $L_{2}=5$. We choose $10^{2}$ 1 as the common node. The common circuits ($C_{[5]}S^{2}$ and $C_{[5]}C_{[2]}S$) have lengths $l_{\rm c1}=5$ and $l_{\rm c2}=7$. There is a single outer cycle ($S$) which has length $l_{\rm o1}=1$. Using these in Theorem 2 yields
$B_{n}=\delta_{n,0}-\delta_{n,3}-\delta_{n,5}+B_{n-1}+B_{n-3}-B_{n-4}+2B_{n-5}-B_{n-6}+B_{n-7}.$ (14)
The generating function for $S_{n}$ is therefore
$\frac{1}{x^{4}}\biggl{(}\frac{1-x^{3}-x^{5}}{1-x-x^{3}+x^{4}-2x^{5}+x^{6}-x^{7}}-\frac{1-x^{4}}{1-x}\biggr{)},$
@@
The generating function $G(x)$ for the number of restricted combinations of $\mathbb{N}_{n}$ with disallowed differences $\{1,5\}$ is given by
$G(x)=\frac{1+2x+2x^{2}+2x^{3}+2x^{4}+4x^{5}+3x^{6}+2x^{7}+x^{8}+x^{9}}{1-x^{2}-x^{3}-x^{4}+x^{5}-x^{6}-x^{7}-x^{8}-x^{10}}.$ (15)
@@
In this case, the $\mathcal{Q}$-comb is a $(2,3,1)$-comb and the metatile-generating digraph (Fig. 1) contains 4 inner cycles of which two are errant loops. Theorem 6 can therefore be applied and we choose the $10^{3}$ 1 node as the pseudo-common node $\mathcal{P}$. The inner cycles passing through $\mathcal{P}$ ($S^{2}C_{[4]}$ and $C_{[2]}SC_{[4]}$) have lengths $L_{1}=4$ and $L_{2}=6$ and their respective errant loops ($C_{[3]}$ and $C_{[2]}$) have lengths $L_{3}=3$ and $L_{4}=2$. The common circuits ($C_{[6]}S^{3}$ and $C_{[6]}C_{[2]}S^{2}$) have lengths $l_{\rm c1}=6$ and $l_{\rm c2}=8$. There is just one outer cycle ($S$) and its length $l_{\rm o1}=1$. From (4) we have
$B_{n}=\delta_{n,0}-\delta_{n,2}-\delta_{n,3}-\delta_{n,4}+\delta_{n,5}+\delta_{n,9}+B_{n-1}+B_{n-2}-2B_{n-5}+2B_{n-6}-B_{n-9}+B_{n-10}-B_{n-11}.$
The generating function for $S_{n}$ is therefore
$G(x)=\frac{1}{x^{5}}\biggl{(}\frac{1-x^{2}-x^{3}-x^{4}+x^{5}+x^{9}}{1-x-x^{2}+2x^{5}-2x^{6}+x^{9}-x^{10}+x^{11}}-\frac{1-x^{5}}{1-x}\biggr{)},$
@@
The proof of the second bijection (Theorem 11) requires the following bijection established in [10] between strongly restricted permutations and tiling with $(\frac{1}{2},g)$-fences (which are just $(\frac{1}{2},g,\frac{1}{2})$-combs) where $g$ is always a nonnegative integer. As the teeth (originally called posts in the context of fences) are not of integer width, it is convenient to regard each cell as being divided into two equal slots into which a tooth can fit [11].
@@ 4 Connections with compositions
OEIS sequences corresponding to a $\mathcal{Q}$ that Theorem 14 applies to along with the set itself are as follows: A00045 {1}, A006498 {2}, A000930 {1,2}, A079972 {2,3}, A003269 {1,2,3}, A351874 {1,3,4}, A121832 {2,3,4}, A003520 {1,2,3,4}, A259278 {2,3,4,5}, A005708 {1,2,3,4,5}, A276106 {2,3,4,5,6}, A005709 {1,2,3,4,5,6}, A322405 {2,3,4,5,6,7}, A005710 {1,2,3,4,5,6,7}, A368244 {2,3,4,5,6,7,8}.
@@
Here the $\mathcal{Q}$-comb is an $(l,g,m,h,r;t)$-comb. The $t=1,2$ cases are covered by the proof of Theorem 14. First note that if $t$ is odd, the length of the comb $q+1=\frac{1}{2}(t-1)(l+g+m+h)+r$ and it is $\frac{1}{2}t(l+g)+(\frac{1}{2}t-1)(m+h)+r$ if $t$ is even from which the respective conditions on $r$ (so that there are a finite number of metatiles) follow. There are two mutually exclusive sets of conditions that ensure that the metatiles contain no more than two combs. Condition (i) that $l\geq g+m+h$ (see Fig. 4(b)) arises as a result of the observation that if $l$ were one less than $g+m+h$ and a second comb were placed starting at the first gap of the first comb, the start a third comb could be placed starting at the final cell of the second gap of the first comb. The first part of condition (ii), namely that $l\geq g,h$ is for the same reasons as the $l\geq g$ condition in Theorem 14. We must also have $l+g+m\geq g+m+h$ (see Fig. 4(c)) which also implies that $l\geq h$. Were $l+g+m$ one less than $g+m+h$, it can be seen that after placing the second comb at the start of the first gap, a third could be placed at the end of the second gap of the first comb. At the same time, we must ensure that $m$ is not too small; otherwise the start of a third comb could be inserted at the start of the second gap of the first comb. Thus $l-1+g\leq m$ (see Fig. 4(d)). The result then follows from Lemma 13. ∎
@@
The following theorem originates from Theorem 6 in [1]. Although there appears to be only one sequence in the OEIS that it applies to (namely, A224809 for which $\mathcal{Q}$ is {2,4}), there are 18 other cases of $\mathcal{Q}$ it applies to for $q\leq 9$ [1].
@@ 5 Connections with bit strings
Others have noted that the number of length-$n$ bit strings (or binary words) having no two 1s that have positions in the string that differ by $q$ (or, equivalently, are separated from one another by $q-1$ digits) is $S^{\{q\}}_{n}$ [15]. This can be generalized by considering the following bit string $s$ representing subset $\mathcal{S}$: the $j$-th bit from the right of $s$ is 1 if and only if $j\in\mathcal{S}$. (This representation is used in the program in the Appendix.) Then it is clear that $S_{n,k}$ is the number of length-$n$ bit strings that contain $k$ 1s placed in such a way that the difference in positions of any two 1s does not equal an element of $\mathcal{Q}$. We now mention an instance of this that corresponds to A130137 in the OEIS [20]. If one instead places the same restrictions on 0s rather than 1s, for the $\mathcal{Q}=\{1\}$ case, the bit strings (i.e., those containing no 00) are called Fibonacci binary words, so named as the number of such words of length $n$ is $F_{n+2}$. A130137 is described as the number of length-$n$ Fibonacci binary words that do not contain the substring 0110. This is $S^{\{1,3\}}_{n}$ since the three other possible disallowed substrings that only prevent the occurrence of two 0s whose positions differ by 3 (i.e., $0$, 0010, and 0100) do not need to be mentioned; they each contain 00 and so cannot be present in Fibonacci binary words.
@@
We show that the $k$-subsets of $\mathbb{N}_{n-q}$ satisfying the conditions for disallowed differences specified by $\mathcal{Q}$ are the same as the sets representing the equivalence classes of binary words of length $n$ in which $\omega$ appears $k$ times. In general, the possible positions of any subword of length $l$ in a word of length $n$ (and therefore the possible elements of the set representing an equivalence class) are $1,\ldots,n-l+1=n-q$. From a restricted subset $\mathcal{S}$ of $\mathbb{N}_{n-q}$ we construct an equivalence class of words as follows. Starting with a length-$n$ word of $x$s (each $x$ representing 0 or 1), for each $s\in\mathcal{S}$ we place the start of subword $\omega$ at position $s$ in the word. Note that $\lfloor\omega/2^{j}\rfloor$ gives the $l-j$ leftmost bits of $\omega$ and $\omega\bmod 2^{l-j}$ gives the $l-j$ rightmost bits. If the two are equal (and so $j\notin\mathcal{Q}$) then, on shifting the digits of subword $\omega$ by $j$ places to the right and overlaying them on the original subword, none of the digits of the original subword are changed. This means that we have constructed a valid equivalence class once we have specified the restrictions on the remaining $x$s to avoid any further appearances of $\omega$. Note that if there an $\omega$ at position $i$ and no other at $i+j$ for all $j=1,\ldots,q$, we may place another $\omega$ at position $i+l$ or after this. This coincides with the fact that $l=q+1$ or any larger integers are not contained in the set $\mathcal{Q}$ of disallowed shifts. The reversible nature of the construction establishes the bijection. ∎
@@
To obtain Corollary 22 to Theorem 20, we require the following result.
Lemma 21 (Corollary 1 in [1]).
If the elements $q_{i}$ of $\mathcal{Q}$ are a well-based sequence then the generating function for $S^{\mathcal{Q}}_{n}$ is given by
$G(x)=\frac{c}{(1-x)c-x},$ (19)
where $c=1+\sum_{i=1}^{\left\lvert\mathcal{Q}\right\rvert}x^{q_{i}}$.
@@
The generating function $\tilde{G}(x)$ for $E^{(\omega)}_{n}$, the number of equivalence classes of length-$n$ binary words with respect to a length-$l$ subword $\omega$ is given by
$\tilde{G}(x)=\frac{\displaystyle{1-\sum_{i=1}^{\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert}\!\!x^{p_{i}}}}{\displaystyle{(1-x)\biggl{(}1+\sum_{i=1}^{\left\lvert\mathcal{Q}\right\rvert}x^{q_{i}}\biggr{)}-x}},$ (20)
where $\mathcal{Q}$ is obtained from $\omega$ using the procedure in Theorem 19, $q_{i}$ are the elements of $\mathcal{Q}$ of which the largest is $q$ (with $q=0$ when $\mathcal{Q}=\{\}$), and $p_{i}$ are the elements of $\mathbb{N}_{q}-\mathcal{Q}$, provided that $q=l-1$.
@@
We first deal with the $q=0$ case. Then $\mathcal{Q}=\{\}$ and the sums in (20) are omitted leaving $\tilde{G}(x)=1/(1-2x)$. In this case, each binary word is in an equivalence class of its own and so $E^{(\omega)}_{n}=2^{n}$ which coincides with the values given by $\tilde{G}(x)$. For $q>0$, since $E^{(\omega)}_{n}=1$ for $10\leq n\leq q-$ 1, and, by Theorem 19, $E^{(\omega)}_{n}=S^{\mathcal{Q}}_{n-q}$ for $n\geq q$, we have, on using Lemma 21, that
$\tilde{G}(x)=\frac{x^{q}c}{(1-x)c-x}+1+x+\cdots+x^{q-1}=\frac{x^{q}c}{(1-x)c-x}+\frac{1-x^{q}}{1-x}.$
This leads to (20) on using the result that $c+\sum_{i=1}^{\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert}x^{p_{i}}=x+x^{2}+\cdots+x^{q}$. ∎
Theorem 5 of [1] also gives a recursion relation for $B^{\mathcal{Q}}_{n,k}$ when the elements of $\mathcal{Q}$ form a well-based sequence. Using Theorems 19 and 1 it is therefore straightforward to obtain a recursion relation for the number of equivalence classes with respect to length-$l$ subword $\omega$ with associated set $\mathcal{Q}$ (satisfying $q=l-1$) of length-$n$ binary words that have a set representation of size $k$.

### v2

@@ 3 Bijections between strongly restricted permutations and restricted combinations
A strongly restricted permutation $\pi$ of the set $\mathbb{N}_{n}$ is a permutation for which the number of permissible values of $\pi(i)-i$ for each $i\in\mathbb{N}_{n}$ is less than a finite number independent of $n$ [16]. Here we give two bijections between restricted combinations and types of strongly restricted permutations where the permutations satisfy $\pi(i)-i\in\mathcal{D}$. Our first bijection concerns a strongly restricted permutation where $\mathcal{D}=\{-1,0,1\}$ and any $m$ consecutive $\pi(i)$ cannot differ from one another by more than $m$.
@@
The permutation $\pi$ of $\mathbb{N}_{n+1}$ corresponding to a subset $\mathcal{S}$ of $\mathbb{N}_{n}$ is formed as follows. If $i\in\mathcal{S}$ (where $i=1,\ldots,n$) then $\pi(i)=i+1$ and $\pi(i+1)=i$ (i.e., a pair of exchanged digits) whereas if neither $i$ nor $i-1$ are in $\mathcal{S}$ (for $i=2,\ldots,n+1$) then $\pi(i)=i$. As $\mathcal{Q}$ contains 1, the pairs of exchanged digits in the permutation cannot overlap. Given that condition (a) is satisfied, the only way for (b) to be violated would be if $\pi(i+1)=i$ and $\pi(i+m)=i+m+1$; this could only occur if both $i$ and $i+m$ were in $\mathcal{S}$ which is impossible since $m\in\mathcal{Q}$. The process of forming a permutation from a subset is clearly reversible and so the bijection is established. ∎
Theorem 8 has connections to the following sequences in the OEIS [20]: A000930 ($S^{\{1,2\}}_{n}$) and A102547 ($S^{\{1,2\}}_{n,k}$), its associated triangle; A130137 ($S^{\{1,3\}}_{n}$), which we mention again in the section on bit strings; A263710 ($S^{\{1,4\}}_{n}$), which was the sequence that led to the discovery of the bijection and we will now prove the conjectured generating function for that entry; and A374737 ($S^{\{1,5\}}_{n}$), which was recently to the OEIS added by the author.
@@
The generating function $g^{\{1,4\}}(x)$ for the number of restricted combinations of $\mathbb{N}_{n}$ with disallowed differences $\{1,4\}$ is given by
$g^{\{1,4\}}(x)=\frac{1+x+x^{2}+x^{3}+2x^{4}+x^{6}}{1-x-x^{3}+x^{4}-2x^{5}+x^{6}-x^{7}}.$ (3)
@@
The $\mathcal{Q}$-comb is a $(2,2,1)$-comb. The metatile-generating digraph for tiling with squares and such combs (Fig. 3) has two inner cycles ($SC_{[3]}$ and $C_{[2]}C_{[3]}$) which have lengths $L_{1}=3$ and $L_{2}=5$. We choose $10^{2}$ 1 as the common node. The common circuits ($C_{[5]}S^{2}$ and $C_{[5]}C_{[2]}S$) have lengths $l_{\rm c1}=5$ and $l_{\rm c2}=7$. There is a single outer cycle ($S$) which has length $l_{\rm o1}=1$. Then from Theorem 5 and Corollary 7,
$g^{\{1,4\}}(x)=\frac{1}{x^{4}}\biggl{(}\frac{1-x^{3}-x^{5}}{1-x^{3}-x^{5}-x+x^{4}+x^{6}-x^{5}-x^{7}}-\frac{1-x^{4}}{1-x}\biggr{)},$
@@
The generating function $g^{\{1,5\}}(x)$ for the number of restricted combinations of $\mathbb{N}_{n}$ with disallowed differences $\{1,5\}$ is given by
$g^{\{1,5\}}(x)=\frac{1+2x+2x^{2}+2x^{3}+2x^{4}+4x^{5}+3x^{6}+2x^{7}+x^{8}+x^{9}}{1-x^{2}-x^{3}-x^{4}+x^{5}-x^{6}-x^{7}-x^{8}-x^{10}}.$ (4)
@@
In this case, the $\mathcal{Q}$-comb is a $(2,3,1)$-comb and the metatile-generating digraph (Fig. 4) falls into the category covered by Theorem 6. We choose the $10^{3}$ 1 node as the pseudo-common node $\mathcal{P}$. There is just one outer cycle ($S$). It is plain and its length $l_{\rm o1}=1$. The common circuits ($C_{[6]}S^{3}$ and $C_{[6]}C_{[2]}S^{2}$) are non-plain and have lengths $\tilde{l}_{\mathrm{c}1}=6$ and $\tilde{l}_{\mathrm{c}2}=8$; their respective errant loops ($C_{[3]}$ and $C_{[2]}$) have lengths $l_{\mathrm{ce}1}=3$ and $l_{\mathrm{ce}2}=2$. The inner cycles passing through $\mathcal{P}$ ($S^{2}C_{[4]}$ and $C_{[2]}SC_{[4]}$) are non-plain and have lengths $\tilde{L}_{1}=4$ and $\tilde{L}_{2}=6$; their respective errant loops ($C_{[3]}$ and $C_{[2]}$) have lengths $l_{\mathrm{e}1}=3$ and $l_{\mathrm{e}2}=2$. From (2) and Corollary 7 we have
$G(x)=\frac{1}{x^{5}}\left(\frac{\displaystyle{1}}{\displaystyle{1-x-\frac{\displaystyle{\frac{\displaystyle{x^{6}}}{\displaystyle{1-x^{3}}}+\frac{\displaystyle{x^{8}}}{\displaystyle{1-x^{2}}}}}{\displaystyle{1-\frac{\displaystyle{x^{4}}}{\displaystyle{1-x^{3}}}-\frac{\displaystyle{x^{6}}}{\displaystyle{1-x^{2}}}}}}}-\frac{1-x^{5}}{1-x}\right),$
@@
The proof we give of the second bijection (Theorem 12) requires the following bijection established in [9] between strongly restricted permutations and tiling with $(\frac{1}{2},g)$-fences (which are just $(\frac{1}{2},g,\frac{1}{2})$-combs) where $g$ is always a nonnegative integer. As the teeth (originally called posts in the context of fences) are not of integer width, it is convenient to regard each cell as being divided into two equal slots into which a tooth can fit [10].
@@ 4 Connections with compositions
OEIS sequences corresponding to a $\mathcal{Q}$ that Theorem 15 applies to along with the set itself are as follows: A000045 {1}, A006498 {2}, A000930 {1,2}, A079972 {2,3}, A003269 {1,2,3}, A351874 {1,3,4}, A121832 {2,3,4}, A003520 {1,2,3,4}, A375985 {1,3,4,5}, A259278 {2,3,4,5}, A005708 {1,2,3,4,5}, A276106 {2,3,4,5,6}, A005709 {1,2,3,4,5,6}, A322405 {2,3,4,5,6,7}, A005710 {1,2,3,4,5,6,7}, A368244 {2,3,4,5,6,7,8}.
@@
Here the $\mathcal{Q}$-comb is an $(l,g,m,h,r;t)$-comb. The $t=1,2$ cases are covered by the proof of Theorem 15. First note that if $t$ is odd, the length of the comb $q+1=\frac{1}{2}(t-1)(l+g+m+h)+r$ and it is $\frac{1}{2}t(l+g)+(\frac{1}{2}t-1)(m+h)+r$ if $t$ is even from which the respective conditions on $r$ (so that there are a finite number of metatiles) follow. There are two mutually exclusive sets of conditions that ensure that the metatiles contain no more than two combs. Condition (i) that $l\geq g+m+h$ (see Fig. 6(b)) arises as a result of the observation that if $l$ were one less than $g+m+h$ and a second comb were placed starting at the first gap of the first comb, the start a third comb could be placed starting at the final cell of the second gap of the first comb. The first part of condition (ii), namely that $l\geq g,h$ is for the same reasons as the $l\geq g$ condition in Theorem 15. We must also have $l+g+m\geq g+m+h$ (see Fig. 6(c)) which also implies that $l\geq h$. Were $l+g+m$ one less than $g+m+h$, it can be seen that after placing the second comb at the start of the first gap, a third comb could be placed at the end of the second gap of the first comb. At the same time, we must ensure that $m$ is not too small; otherwise the start of a third comb could be inserted at the start of the second gap of the first comb. Thus $l-1+g\leq m$ (see Fig. 6(d)). The result then follows from Lemma 14. ∎
@@
The following theorem originates from Theorem 6 in [2]. The OEIS sequences to which it applies along with the corresponding $\mathcal{Q}$ are as follows: A224809 {2,4}, A375981 {1,4,5}, A375982 {2,3,5}, A375983 {2,4,5}.
@@ 5 Connections with bit strings
Others have noted that the number of length-$n$ bit strings (or binary words) having no two 1s that have positions in the string that differ by $q$ (or, equivalently, are separated from one another by $q-1$ digits) is $S^{\{q\}}_{n}$ [14]. This can be generalized by considering the following bit string $s$ representing subset $\mathcal{S}$: the $j$-th bit from the right of $s$ is 1 if and only if $j\in\mathcal{S}$. (This representation is used in the program in the Appendix.) Then it is clear that $S^{\mathcal{Q}}_{n,k}$ is the number of length-$n$ bit strings that contain $k$ 1s placed in such a way that the difference in positions of any two 1s does not equal an element of $\mathcal{Q}$. We now mention an instance of this that corresponds to A130137 in the OEIS [20]. If one instead places the same restrictions on 0s rather than 1s, for the $\mathcal{Q}=\{1\}$ case, the bit strings (i.e., those containing no 00) are called Fibonacci binary words, so named as the number of such words of length $n$ is $F_{n+2}$. A130137 is described as the number of length-$n$ Fibonacci binary words that do not contain the substring 0110. This is $S^{\{1,3\}}_{n}$ since the three other possible disallowed substrings that only prevent the occurrence of two 0s whose positions differ by 3 (i.e., $0$, 0010, and 0100) do not need to be mentioned; they each contain 00 and so cannot be present in Fibonacci binary words.
@@
We show that the $k$-subsets of $\mathbb{N}_{n-q}$ satisfying the conditions for disallowed differences specified by $\mathcal{Q}$ are the same as the sets representing the equivalence classes of binary words of length $n$ in which $\omega$ appears $k$ times. In general, the possible positions of any subword of length $l$ in a word of length $n$ (and therefore the possible elements of the set representing an equivalence class) are $1,\ldots,n-l+1=n-q$. From a restricted subset $\mathcal{S}$ of $\mathbb{N}_{n-q}$ we construct an equivalence class of words as follows. Starting with a length-$n$ word of $x$s (each $x$ representing 0 or 1), for each $s\in\mathcal{S}$ we place the start of subword $\omega$ at position $s$ in the word. Note that $\lfloor\omega/2^{j}\rfloor$ gives the $l-j$ leftmost bits of $\omega$ and $\omega\bmod 2^{l-j}$ gives the $l-j$ rightmost bits. If the two are equal (and so $j\notin\mathcal{Q}$) then, on shifting the digits of subword $\omega$ by $j$ places to the right and overlaying them on the original subword, none of the digits of the original subword are changed. This means that we have constructed a valid equivalence class once we have specified the restrictions on the remaining $x$s to avoid any further appearances of $\omega$. Note that if there is an $\omega$ at position $i$ and no other at $i+j$ for all $j=1,\ldots,q$, we may place another $\omega$ at position $i+l$ or after this. This coincides with the fact that $l=q+1$ or any larger integers are not contained in the set $\mathcal{Q}$ of disallowed shifts. The reversible nature of the construction establishes the bijection. ∎
@@
The generating function $\tilde{g}(x,y)$, whose coefficient of $x^{n}y^{k}$ equals $E^{(\omega)}_{n,k}$, the number of equivalence classes of length-$n$ binary words with respect to a length-$l$ subword $\omega$ that have a set representation of size $k$, is given by
$\tilde{g}(x,y)=\frac{1-\bar{c}}{(1-x)c-xy},$ (8)
where $\bar{c}=\sum_{i=1}^{a}x^{p_{i}}y$, $a=\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert$, $c=1+\sum_{i=1}^{\left\lvert\mathcal{Q}\right\rvert}x^{q_{i}}y$, the set $\mathcal{Q}$ is obtained from $\omega$ using the procedure in Theorem 20, $q_{i}$ are the elements of $\mathcal{Q}$ of which the largest is $q$ (with $q=0$ when $\mathcal{Q}=\{\}$), and $p_{i}$ are the elements of $\mathbb{N}_{q}-\mathcal{Q}$, provided that $q=l-1$.
Figure 9: Digraph for tiling a board with squares and $\mathcal{Q}$-combs when the $q_{i}\in\mathcal{Q}$ are a well-based sequence. Node label $\sigma_{i}$ (for $i=1,\ldots,a$) denotes the bit string corresponding to filling the first $i-1$ empty cells of the $\mathcal{Q}$-comb with squares and discarding the leading 1s [2].
@@
We first deal with the $q=0$ case. Then $\mathcal{Q}=\{\}$ and the sums in (8) are omitted leaving $\tilde{g}(x,y)=1/(1-x-xy)$. In this case, each binary word is in an equivalence class of its own and so $E^{(\omega)}_{n,k}=\tbinom{n}{k}$ since this is the number of length-$n$ binary words that contain $k$ 1s. The generating function for $\tbinom{n}{k}$ is $1/(1-x-xy)$. For $q>0$, since $E^{(\omega)}_{n}=E^{(\omega)}_{n,0}=1$ for $10\leq n\leq q-$ 1, and, by Theorem 20, $E^{(\omega)}_{n,k}=S^{\mathcal{Q}}_{n-q,k}$ for $n\geq q$ and from Theorem 1 $S^{\mathcal{Q}}_{n-q,k}=B^{\mathcal{Q}}_{n,k}$ for $n\geq q$, overall we have $E^{(\omega)}_{n,k}=B^{\mathcal{Q}}_{n,k}$. Hence $\tilde{g}(x,y)=G(x,y)$, the generating function such that the coefficient of $x^{n}y^{k}$ is the number of tilings of an $n$ board that contain $k$ combs. The digraph for tiling a board with squares and $\mathcal{Q}$-combs when the elements of $\mathcal{Q}$ are a well-based sequence is shown in Fig. 9 [2]. The digraph has a common node ($\sigma_{1}$) and the inner cycles ($S^{i-1}C_{[p_{i}]}$ for $i=1,\ldots,a$) have lengths $p_{i}$. There is a single common circuit ($C_{[q+1]}S^{a}$), which is of length $q+1$, and one outer cycle ($S_{[1]}$), which is of length 1. The common circuit and all the inner cycles contain a single comb. From Theorem 5 we then have
$G(x,y)=\frac{1-\bar{c}}{1-\bar{c}-x+x\bar{c}-x^{q+1}y}.$ (9)
Using the result that $c+\bar{c}=1+\sum_{i=1}^{q}x^{i}y=1+(x-x^{q+1})y/(1-x)$ to simplify the denominator then gives (8). ∎
The following corollary, which is a generalization of a result obtained previously (Corollary 1 in [2]; Theorem 2 in [12]), is then easily obtained from (9) by applying Corollary 7.
Corollary 23.
If the elements $q_{i}$ of $\mathcal{Q}$ are a well-based sequence then the generating function such that the $x^{n}y^{k}$ coefficient thereof is $S^{\mathcal{Q}}_{n,k}$ is given by
$g(x,y)=\frac{c}{(1-x)c-xy},$ (10)
where $c=1+\sum_{i=1}^{\left\lvert\mathcal{Q}\right\rvert}x^{q_{i}}y$.